# A Tale of Two Imperatives: Privacy and Explainability

### v3

@@ 6 Key Observation
In this section, we present our experimental results to find out whether the post-hoc explainers agree with the Localization Assumption (LA), defined in Section 4.2. Integrated Gradients and Grad-Shap showed more than 45% DS everywhere which directly violates our DS threshold of $15\%$. As we consider DS a sanity check for LA, and these two explainers fail this test, we exclude them from further discussions. Across models, around $30-40\%$ Grad-Cam explanations of test samples violated the DS threshold, we have considered the rest of the samples to calculate PIS. Almost all test samples passed the DS test across models for Saliency and SmoothGrad.
We report our findings for the three networks: DenseNet-121, ResNet-34 and EfficientNet-v2 in Figure 1, 2 and 3 respectively. For each of these networks, we report class-wise results against each of the three explainers which pass the DS test. Against each $\epsilon$, we report average of the PIS values over the test set ($PIS_{Avg}$), $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$, and agreement on the “hard predictions” obtained from $\mathcal{M}$ and $\mathcal{M^{\prime}}$ over the test set ($Agreement$). $PIS_{Avg}$ gives the extent to which an explainer agrees with the LA in a particular setup; $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$ quantifies the Perf Comp. requirement, and $Agreement$ quantifies the Alignment requirement for the private model to be functionally comparable w.r.t. the original model, defined in Section 4.2. We observe all of the explainers across $\epsilon$, classes, and models are performing poorly in terms of $PIS_{Avg}$ [^12] One may argue that what if we measure similarity without segregating attributes? This although disregards our philosophy on similarly measure still we did and to be precise, never found $PIS_{Avg}>0.16$.. In fact, the $PIS_{Avg}$ never crosses $0.32$; such a low $PIS_{Avg}$ indicates that none of the explainers agree with the LA. Almost all cases except for the Normal class against low $\epsilon$ values, $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}\approx 1$, which indicates that almost all private models achieved near equal accuracy as of their base counterparts. Nearly $70\%$ of cases achieved $Agreement>0.50$. There is no apparent relation or trend the $PIS_{Avg}$ follows with $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$ and $Agreement$. All in all, the explanation quality is consistently poor no matter what happens to the other parameters. Empirically, it is clear that all gradient-based explainers disregard LA, but why is it so? We’ll try to find out in the next section.
@@ 7 Why Explainers Performed so Poorly?
As noted in the previous section, the explanations from private and non-private models are largely uncorrelated. Private models are trained using DP-SGD, which involves gradient clipping and adding calibrated noise during backpropagation to achieve DP. To understand the divergent behavior of explainers between these models, we first analyze how DP training alters model parameters. We conduct a comparative study on the (layerwise) representations learned by both private and non-private models, followed by a righrous investigation into the (local) sensitivity of these representations. While the former reveals how DP training modifies the parameter space, the latter explains how these modifications influence the model’s reaction to inputs, ultimately affecting explainer behavior. To the best of our knowledge, this is the first study to comprehensively examine the representations learned by DP models and compare them to non-private models. Moreover, the layer-wise sensitivity analysis has not been previously explored for any model, including DP models. We include the implementation details in Appendix C and will publicly share the weights and codes upon publication.
@@
Statistical tests for activations are often non-trivial and subject to a few constraints such as invariance to permutation of the neurons, orthogonal transformation (Klabunde et al., 2024) etc. Consequently, we use the Hilbert-Schmidt Independence Criterion (HSIC) (Gretton et al., 2007) which operates over a Reproducing Kernel Hilbert Space (RKHS) and can detect dependencies across complex, high-dimensional variables while respecting the aforesaid constraints. We employ the HSIC unconditional independence test using the two-parameter $\gamma$ approximation scheme with a p-value cut-off of $0.05$ (Gretton et al., 2012).
From our findings, we were able to reject the null hypothesis $\mathcal{H}_{0}$ for almost all layers [^13] For example, across models we never crossed $7-10\%$ of the layers for any $\epsilon$ where we accept $\mathcal{H}_{0}$. In other words, across models layerwise representations are not independent, for all $\epsilon$. Since the representations are sufficiently comparable, we’ll now compute their similarity layerwise. However, before a full-fledged comparison, we present a short note on representational similarity (RS).
@@ 7.0.1 On Representational Similarity
Representational similarity measures compare neural networks by computing similarity between activations of a fixed set of inputs (in our case, test set) at a given pair of layers. There’s been an extensive line of work on this domain however, each of them can be catagorized depending on the notion of similarity it regards. Each similarity measure follows a set of assumptions. In our paper, we have considered the most regarded set of assumptions and its corresponding similarity measure, namely Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Gretton et al., 2007). CKA is built upon HSIC and bounded within the (closed) interval of $[0,1]$. However, there are a few (discovered) shortcomings of CKA, as discussed below.
The wildly used CKA with the linear kernel, is equivalent to the RV coefficient and it is already shown that the same seldom yields values close to $1$ due to one of its inherent shortcomings: although the coefficient is inherently constrained to values between 0 and 1, it rarely reaches values near 1 because the denominator is typically much larger relative to its theoretical maximum value (Puccetti, 2022). Also, the lesser adapted RBF-CKA is highly sensitive to the kernel width and ineffective for small values (Davari et al., 2022).
Furthermore, very recently Cui et al. (Cui et al., 2022) have discovered that the inter-example (dis)similarity in the representation space works as a confounder. In their words, “This leads to spuriously high CKAs even between two random neural networks, and counter-intuitive conclusions when comparing CKAs on sets of models trained on different domains ...”. They fixed this problem by regressing out the confounder from the similarity matrices of two representations. Our analysis uses this deconfounded version of CKA (dCKA). For HSIC, we used both linear and non-linear (RBF) kernels but obtained similar results.
@@
All in all, the private models don’t perceive the data in a strongly correlated manner w.r.t their non-private counterparts. However, we argue that this is not the only reason for the explanations to be altered, as explanations are based on the sensitivity of the features subject to a class of interest the model shows. In the second phase of our investigation, we shall look into how much, for an obtained output, the model is sensitive to all the layerwise representations it obtained in the forward pass.
@@ 7.1 Do the models show similar sensitivity?
Due to these inherent challenges, we only conduct hypothesis testing using HSIC, as previously mentioned, to check whether the sensitivity of representations is statistically (in)dependent, where the corresponding outputs match [^16] Essentially, this is the dataset considered for PIS.. Here also, we employed both linear and non-linear (RBF) kernels but obtained similar results.
From our findings, we were unable to reject the null hypothesis $\mathcal{H}_{0}$ for across almost all layers [^17] In no model did we reject $\mathcal{H}_{0}$ for more than 3-7% of the layers across $\epsilon$. Notably, for DenseNet-121, all layers were found to be independent. In other words, unlike representations, the sensitivity of representations is independent across models. Regarding Grad-CAM, we typically use the last layer to generate the CAM. Our experiment showed that the gradients of these layers exhibit independence, meaning the CAMs produced by different models will not be comparable. Similarly, other explainers relying on $\nabla_{\sigma_{l}}\Theta$ cannot produce aligned and substantially comparable explanations, as representations across levels along with the last layer show independent sensitivity between non-private and their private counterparts [^18] A natural question arises: could an explainer, despite of consistently incomparable sets of $\nabla_{\sigma_{l}}\Theta$, produce similar explanations? Based on our experiments, we did not identify any commonly used explainer that exhibited this property for the models selected. Furthermore, we argue the faithfulness of such explainers, if they exist, would be questionable. Conversely, assessing the quality of explanations when (local) sensitivity is consistently comparable across any pair of models falls outside the scope of this paper. .
@@ 8 Is there any alternative way to have private explanations?
We now know that the DP models cannot accommodate widely used post-hoc explainers primarily due to their own behaviour. As a result, we cannot move forward with DP models to obtain private explanations. However, just for private explanations, we don’t need that. We can achieve the same locally as well.
In this approach, rather than training the model with DP-SGD, we use the non-private model as usual and take their explanations. We add calibrated noise to the explanation to make it Local Differential Private (LDP)
@@
Where $\mathcal{A}$ satisfies $\epsilon$-differential privacy for $t=\frac{\Delta f}{\epsilon}$, and $\Delta f$ is the sensitivity [^19] This sensitivity is different from the (local) sensitivity we discussed in section 7. of $f$ Dwork et al. (2014).
Here, Noise is directly applied to the heatmap obtained from the explainer. Following the guidelines by Fan (2018), for an image with $c$ channels, each of which spans $k$ possible pixel intensities, the sensitivity of the query is $(k-1)nc$, where $n$ is the maximum number of pixels that can differ in two adjacent images. Given this query sensitivity, $\epsilon$-differential privacy can be provided by independently applying the Laplace mechanism to each pixel in each channel, using a scaling parameter of $(k-1)nc/\epsilon$. However, such a large amount of noise, especially for smaller $\epsilon$, would completely obfuscate the semantics of the image. To reduce the query sensitivity, the authors propose dividing the image into $b\times b$ grids of pixels, where each grid is averaged to create a uniform value. Treating each grid as a single unit to be obfuscated reduces the required scaling parameter to $(k-1)nc/b^{2}\epsilon$. Therefore, in the case of explanations (or likewise any heatmap with 1 channel), the query sensitivity is $255n/b^{2}$.
Any valid set of $n$ and $b$ chosen by the user ensures $\epsilon$-differential privacy, but tuning these parameters offers a trade-off between privacy and utility. Therefore, this framework allows us to generate explanations that are both useful and private.
Here, a natural question arises: what constitutes a good value for $n$? Previous works have tackle this problem empirically as this primarily depends on the use-case. We, therefore, empirically tested with a diverse set of $(n,b)$ and communicated the results with the physicians of SRM University AP. We get the best response for the tuple (16, 14), and the most competent explainers were Grad-Shap and Integrated Gradients. In our case, we get useful results $\epsilon=4$ onwards. We selected the ResNet-34 for demonstrating our results and we have reported explanations from all four explainers in Appendix A.
@@ 8.1 What Could Be the Notion of PIS and LA Here?
After LDP-fying with $b\times b$ grids, we cannot compare explanations pixel-to-pixel. Here, our motivation is to compare how much LDP-fying degrades the quality of the explanation, to measure that we have to change the similarity measure accordingly. That is why here, unlike global DP, we do not measure DS or correlation but following the postulate of LA, we expect that the most competent explanation should be least affected and thus should be most similar and can be a close proxy to the actual explanation. For the notion of PIS, we measure the similarity of the newly made explanation image with Structural Similarity Index (SSIM) Wang et al. (2004). SSIM is suitable here because it captures perceptual similarities rather than exact pixel values, which is crucial for understanding qualitative degradation. While pixel-by-pixel comparison may not be sensitive to perceptual changes, SSIM evaluates structural components—luminance, contrast, and structure—that reflect images are visually processed (by the observer). In this context, SSIM effectively measures the similarity by capturing how much essential information is retained after privacy transformations, aligning with our goal of assessing how closely the modified explanation approximates the original. The most robust explanation will remain minimally affected, thus showing the highest SSIM, making it a strong proxy for maintaining interpretability under privacy constraints.
However, in this setup, we face a challenge to work with Grad-CAM. It produces a region of interest (ROI) rather than a per-pixel attribute score which is substantially small (a grid of $7\times 7$ or $8\times 8$). Fixing the hyperparameters in the ROI to make it LDP, we found substantially challenging as even minor noise overwhelms the attributions due to the relatively small number of pixels in ROIs, as a result, we have to discard the same from further evaluations. For demonstrating purposes, we have chosen the ResNet-34 for our software and reported the mean SSIM value for each class and explanations in table 8.1. However, here in this setup, we do not put any presuppositions on the baseline unlike global DP and adopt ad-hoc strategy to set the threshold as long as the private explanations are approved by the concerned physician(s) as there is an inherent trade-off between explanations’ quality and privacy guarantee.
Quality of the explanation (SSIM here) evaluates the perceptual similarity of two images by comparing structural and luminance components. However, introducing LDP noise degrades these structural details to protect individual pixel-level information, thereby reducing SSIM. Balancing the degree of noise to achieve a desired privacy level inevitably diminishes SSIM, as high privacy requires more noise, which more significantly impacts the structural fidelity of the image. Our set of hyperparameters craft-fully stricks the balance and produces succinctly practitioner-approved private explanations. In our experiment, we roughly get PIS between $0.4-0.5$ for all explainers.
$\begin{array}[]{c|c|c|c|c}&\text{Grad-SHAP}&\text{Integrated Gradient}&\text{Saliency}&\text{SmoothGrad}\\ \hline\cr\text{Tuberculosis}&0.52&0.50&0.49&0.52\\ \hline\cr\text{Pneumonia}&0.53&0.51&0.43&0.49\\ \hline\cr\text{Normal}&0.56&0.54&0.49&0.53\\ \end{array}$
In this setup, the model is non-private but the explanations are private, we name this setup as Hybrid DP. Our novel software is outlined using this setup in Figure 9.
@@
9 The Novel Software
Based on our extensive findings, we propose a comprehensive software architecture that ensures both privacy preservation and model explainability. The system is designed with a scrutiny check and privacy-preserving components, orchestrated to work seamlessly while maintaining both RTP and RTE guarantees.
System Overview and Security Measures: At the entry point, all incoming images pass through an autoencoder (AE) that acts as the first line of security Neloy and Turgeon (2024); Berahmand et al. (2024). We get $94.7\%$ mean accuracy over anomalous inputs202020we, after consulting with physicians from SRM University AP, have considered images from classes other than Pneumonia and TB such as Cardiomegaly, Aortic enlargement here as anomalous inputs; we have randomly taken 50 samples per class from Nguyen et al. (2022) dataset. at a reconstruction loss threshold ($\kappa$) of $0.04$. This AE serves dual purposes: (i) it validates the authenticity of inputs by detecting anomalous patterns that might indicate adversarial attacks or corrupted data, and (ii) it ensures that only legitimate medical images enter the system’s processing pipeline. Images that deviate from $\kappa$ are rejected, preventing potential security vulnerabilities.
Core Processing Pipeline: The validated images then enter the main processing pipeline, where our non-private model resides at the center of the system. This model is secured behind an API interface. It generates the prediction for given images which, over an encrypted channel, is received by the concerned physician(s). Note to mention in this setup, we do not make the output private intentionally as adding noise in the softmax vector may deceive the physician if the magnitude of noise is too high. Next, it also causes a drop in utility. Finally, as we only communicate the final label and the LDP-fied explanation with the patients (subject to physicians’ approval), we do not need to disclose this information and we securely communicate the same over an encrypted channel with the concerned physician.
Generating Privacy Preserving Explanation: The system then proceeds to the explanation generation phase, where post-hoc explainers analyze the model’s decision. These explanations undergo a critical privacy-preserving transformation in what we term the noise chamber. Here, carefully calibrated noise is injected into the explanations to make the generated explanations LDP. The noise calibration is based on our theoretical findings about sensitivity patterns in DP models, ensuring that the added noise preserves the essential aspects of the explanations while providing privacy guarantees.
Quality Assesment and Information Delivery: After noise injection, the system performs an SSIM analysis between the LDP-fied explanations and their original counterparts. This quality check ensures that the privacy-preserving mechanisms haven’t significantly degraded the explanation quality. The system prioritizes the top-K (for our case, $K=1-2$) explanations. This is a crucial step as explanations which are severely degraded (say SSIM is in negative) should be discarded to save the time and effort of the concerned physician(s). These selected explanations, along with the model’s predictions, over the encrypted channels, are securely communicated to the concerned physician(s).
Final Communication Protocol: Once the concerned physician(s) formally reviews the prediction and explanations, we communicate the same to the patients subject to concerned physician(s)’ approval.

### v4

@@ 6 Key Observation
In this section, we present our experimental results to find out whether the post-hoc explainers agree with the Localization Assumption (LA), defined in Section 4.2. For Integrated Gradients and Grad-Shap, across all models and test samples chosen, we obtained DS of more than 45%, which directly violates our DS threshold of $15\%$. As we consider DS as a sanity check for LA, and these two explainers fail this test, we exclude them from further discussions. Across models, for around $30-40\%$ of all test samples, Grad-Cam explanations violated the DS threshold of $15\%$; consequently, we have considered the rest of the test samples to calculate PIS for Grad-Cam. All test samples passed the DS test across models for Saliency and SmoothGrad.
We report our findings for the three networks: DenseNet-121, ResNet-34, and EfficientNet-v2 in Figure 2, 2, and 3 respectively. We report class-wise results for each of these networks against each of the three explainers that pass the DS test. Against each $\epsilon$, we report average of the PIS values over the test set ($PIS_{Avg}$), $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$, and agreement on the “hard predictions” obtained from $\mathcal{M}$ and $\mathcal{M^{\prime}}$ over the test set ($Agreement$). $PIS_{Avg}$ gives the extent to which an explainer agrees with the LA in a particular setup; $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$ quantifies the Perf Comp. requirement, and $Agreement$ quantifies the Alignment requirement for the private model to be functionally comparable w.r.t. the non-private counterpart, defined in Section 4.2. We observe all of the explainers across $\epsilon$, classes, and models are performing very poorly in terms of $PIS_{Avg}$ [^6] One may argue that what if we measure similarity without segregating attributes? This, although disregards our philosophy on similarly measure, still we did, and to be precise, never found $PIS_{Avg}>0.16$.. In fact, the $PIS_{Avg}$ never crosses $0.32$; such a low $PIS_{Avg}$ indicates that none of the explainers agree with the LA. In almost all cases, except for the Normal class against low $\epsilon$ values, $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}\approx 1$, which indicates that almost all private models achieved near equal accuracy as of their base counterparts. Nearly $70\%$ of cases achieved $Agreement>0.50$. There is no apparent relation or trend the $PIS_{Avg}$ follows with $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$ and $Agreement$. All in all, the explanation quality is consistently poor, no matter what happens to the other parameters. We observe the very same trend for another benchmark dataset we considered (see Appendix Appendix A.). Empirically, it is clear that all gradient-based explainers disregard LA and thus seem not to be suitable with DP models. With that note, we (partially) answered our primary research question that DP models and off-the-shelf explainers (apparently) don’t go well together. But why is it so? We’ll try to find out in the next section. Also, we’ll explore an alternative route to generate private explanations in the subsequent section.
@@ 7 Why Explainers Performed so Poorly?
As noted in the previous section, the explanations from private and non-private models are largely uncorrelated. Private models are trained using DP-SGD, which involves gradient clipping and adding calibrated noise during backpropagation to achieve DP. To understand the divergent behavior of explainers between these models, we first analyze how DP training alters model parameters. We conduct a comparative study on the (layerwise) representations learned by both private and non-private models, followed by a rigorous investigation into the (local) sensitivity of these representations. While the former reveals how DP training modifies the parameter space, the latter explains how these modifications influence the model’s reaction to inputs, ultimately affecting explainer behavior. To the best of our knowledge, this is the first study to comprehensively examine the representations learned by DP models and compare them to non-private models. Moreover, the layer-wise sensitivity analysis has not been previously explored for any model, including DP models. We include the implementation details in Appendix C.
7.1 Did the models perceive the same information?
@@
Statistical tests for activations are often non-trivial and subject to a few constrain,ts such as invariance to permutation of the neurons, orthogonal transformation (Klabunde et al., 2024) etc. Consequently, we use the Hilbert-Schmidt Independence Criterion (HSIC) (Gretton et al., 2007) which operates over a Reproducing Kernel Hilbert Space (RKHS) and can detect dependencies across complex, high-dimensional variables while respecting the aforesaid constraints. We employ the HSIC unconditional independence test using the two-parameter $\gamma$ approximation scheme with a p-value cut-off of $0.05$ (Gretton et al., 2012). For HSIC, we used both linear and non-linear (RBF) kernels, but obtained similar results.
From our findings, we were able to reject the null hypothesis $\mathcal{H}_{0}$ for almost all layers [^7] For example, across models we never crossed $7-10\%$ of the layers for any $\epsilon$ where we accept $\mathcal{H}_{0}$. In other words, across models, layerwise representations are not independent, for all $\epsilon$. Since the representations are sufficiently comparable, we’ll now compute their similarity layerwise. However, before a full-fledged comparison, we present a short note on representational similarity (RS).
@@ 7.1.1 On Representational Similarity
Representational similarity measures compare neural networks by computing similarity between activations of a fixed set of inputs (in our case, the test set) at a given pair of layers. There’s been an extensive line of work on this domain however, each of them can be categorized depending on the notion of similarity it regards. Each similarity measure follows a set of assumptions. In our paper, we have considered the most regarded set of assumptions and its corresponding similarity measure, namely Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Gretton et al., 2007). CKA is built upon HSIC and bounded within the (closed) interval of $[0,1]$. However, there are a few (discovered) shortcomings of CKA, as discussed below.
The wildly used CKA with the linear kernel, is equivalent to the RV coefficient and it is already shown that the same seldom yields values close to $1$ due to one of its inherent shortcomings: although the coefficient is inherently constrained to values between 0 and 1, it rarely reaches values near 1 because the denominator is typically much larger relative to its theoretical maximum value (Puccetti, 2022). Also, the less adapted RBF-CKA is highly sensitive to the kernel width and ineffective for small values (Davari et al., 2022).
Furthermore, very recently, Cui et al. (Cui et al., 2022) have discovered that the inter-example (dis)similarity in the representation space works as a confounder. In their words, “This leads to spuriously high CKAs even between two random neural networks, and counter-intuitive conclusions when comparing CKAs on sets of models trained on different domains ...". They fixed this problem by regressing out the confounder from the similarity matrices of two representations. Our analysis uses this deconfounded version of CKA (dCKA).
@@
All in all, the private models don’t perceive the data in a strongly correlated manner w.r.t their non-private counterparts. However, we argue that this is not the only reason for the explanations to be altered, as explanations are based on the sensitivity of the features subject to a class of interest, the model shows. In the second phase of our investigation, we shall look into how much, for an obtained output, the model is sensitive to all the layerwise representations it obtained in the forward pass.
@@ 7.2 Do the models show similar sensitivity?
Due to these inherent challenges, we only conduct hypothesis testing using HSIC, as previously mentioned, to check whether the sensitivity of representations is statistically (in)dependent, where the corresponding outputs match [^10] Note, this is the dataset considered for PIS.. Here also, we employed both linear and non-linear (RBF) kernels but obtained similar results.
From our findings, we were unable to reject the null hypothesis $\mathcal{H}_{0}$ for almost all layers [^11] In no model did we reject $\mathcal{H}_{0}$ for more than 3-7% of the layers across $\epsilon$. Notably, for DenseNet-121, all layers were found to be independent. In other words, unlike representations, the sensitivity of representations is independent across models. Regarding Grad-CAM, we typically use the last layer to generate the CAM. Our experiment showed that the gradients of these layers exhibit independence, meaning the CAMs produced by different models will not be comparable. Similarly, other explainers relying on $\nabla_{\sigma_{l}}\Theta$ cannot produce aligned and substantially comparable explanations, as representations across levels along with the last layer show independent sensitivity between non-private and their private counterparts [^12] A natural question arises: could an explainer, despite of consistently incomparable sets of $\nabla_{\sigma_{l}}\Theta$, produce similar explanations? Based on our experiments, we did not identify any commonly used explainer that exhibited this property for the models selected. Furthermore, we argue on the faithfulness of such explainers, if they exist, would be questionable. Conversely, assessing the quality of explanations when (local) sensitivity is consistently comparable across any pair of models falls outside the scope of this paper. .
@@ 8 Is there any alternative way to have private explanations?
We now know that the DP models cannot accommodate widely used post-hoc explainers primarily due to their own nature. As a result, we cannot move forward with DP models to obtain private explanations that will be useful as a proxy to that of the non-DP counterpart. However, just for private explanations, we don’t need that. We opt for an alternative route to achieve the same locally.
In this approach, rather than training the model with DP-SGD, we use the non-private model as usual and take its explanations. We add calibrated noise to the explanation to make it Local Differential Private (LDP)
@@
Where $\mathcal{A}$ satisfies $\epsilon$-differential privacy for $t=\frac{\Delta f}{\epsilon}$, and $\Delta f$ is the sensitivity [^13] This sensitivity is different from the (local) sensitivity we discussed in section 7. of $f$ (Dwork et al., 2014).
Here, noise is directly added to the heatmap (explanation) obtained from the explainer. Following Fan (Fan, 2018), the sensitivity of a query for an image with $c$ channels and $k$ possible pixel intensities is $(k-1)nc$, where $n$ is the maximum number of differing pixels. To ensure $\epsilon$-differential privacy, if we apply the Laplace mechanism to each pixel in each channel, scaling by $(k-1)nc/\epsilon$, such noise can obscure image semantics, particularly for smaller $\epsilon$. To mitigate this, Fan (Fan, 2018) advises dividing the image into $b\times b$ grids, averaging each grid’s pixels, which reduces the sensitivity ($\Delta f$) to $(k-1)nc/b^{2}$. Therefore, for images (including heatmaps) with a single channel, the sensitivity becomes $255n/b^{2}$. We empirically tested with a diverse set of $(n,b)$ for LDP explanations and communicated the results with the concerned physicians. We get the best response for the tuple ($16,14$), and the most competent explainers were Grad-Shap and Integrated Gradients. In our case, we get useful results $\epsilon=4$ onwards. We selected the ResNet-34 to demonstrate our results for all four explainers, which can be found in the supplementary material.
@@ 8.1 What Could Be the Notion of PIS and LA Here?
After LDP-fying with $b\times b$ grids, pixel-wise comparisons are impractical. However, the postulates of LA does hold here as well: the most competent explanations should be least affected by LDP transformations. Therefore, such private explanations ($s^{\prime}$) should be most similar to their non-private counterpart (s) and serve as a close proxy. To quantify the degradation of the quality of $s^{\prime}$ LDP, we employ the Structural Similarity Index (SSIM) (Wang et al., 2004) as PIS in this context. In the absence of pixel-by-pixel comparison, unlike global DP, SSIM aims to reflect how much essential information is retained post-LDPfication, primarily focusing on several perceptual similarities by evaluating structural components in $s^{\prime}$. However, as balancing the degree of noise to achieve a desired privacy level may severely affect the structural fidelity of $s^{\prime}$, thus inevitably diminishes SSIM, we primarily use SSIM evaluation as an elimination test to discard any obfuscated LDP-fied explanations (having substantially low, near zero, or negative SSIM). Furthermore, we eliminated Grad-CAM as its small region of interests ($7\times 7$ or $8\times 8$) tend to get overwhelmed by LDP noise, and fixing the new set of hyperparameters here is also challenging. We roughly get PIS between $0.4-0.5$ for all other explainers (Table 2); however, the concerned physician recommended Integrated Gradients and Grad-Shap to be most useful throughout our experiments.
Grad-SHAP Integrated Gradient Saliency SmoothGrad
Tuberculosis 0.52 0.50 0.49 0.52
Pneumonia 0.53 0.51 0.43 0.49
Normal 0.56 0.54 0.49 0.53
Table 2: Mean SSIM for LDP-fied explanations.
In this setup, the model is non-private, but the explanations are private; we name this setup as Hybrid DP. Our novel software is outlined using this setup in Figure 7.
@@
Table 3: Comparison of Privacy Methods
9 The Novel Software Pipeline
So far, we have investigated non-trivial intricacies of DP models and XAI methods, and we figured out an alternative way to generate private explanations using LDP. Based on all the insights we have consolidated so far, now we propose a privacy-preserving software pipeline that aims to reconcile model explainability and privacy [^14] This proposed Software Pipeline is exclusively based upon our exhaustive findings and insights that we have elaborated throughout our study, we do not claim this to be the (only) ideal setup for high-stakes..
System Overview and Security Measures: At the system’s entry point, all incoming medical images pass through a trained autoencoder (AE) (Neloy and Turgeon, 2024), which filters out anomalous inputs. We achieve $94\%$ mean accuracy on anomaly detection at a reconstruction loss threshold ($\kappa=0.07$), using randomly sampled 50 test images from each non-target class (Cardiomegaly, Aortic enlargement from (Nguyen et al., 2022) dataset) as anomalous examples. Anomalous sample classification can be done with various methods, but we chose AE due to its widespread use in anomaly detection (Neloy and Turgeon, 2024).
Core Processing Pipeline: Validated images go through the core processing pipeline, where a non-private model makes predictions. The model’s prediction (only the label) and the LDP-fied explanation (described in Section 8) are generated. These explanations are assessed for quality using SSIM (as shown in Figure 7). Explanations with poor SSIM scores (e.g., negative or near-zero values) are discarded to avoid wasting the physician’s time.
The remaining top-K highest-quality LDP-fied explanations (here, K = 1-2), along with the model’s prediction, are securely transmitted to the physician via an encrypted channel. This information is sent to the physician(s), not the patient, at this moment.
Final Communication Protocol: The final prediction and explanation, upon physicians’ approval, are communicated to the patient.
$\blacksquare$ How much ‘Compromise’ is required to ‘Reconcile’? In our proposed software pipeline, we refute DP models for reconciling privacy and explainability in high-stakes, as XAI methods and DP models don’t go together. Therefore, we must clearly compromise as we have to fall back to non-private model when we aim to gracefully balance both privacy and explainability in high-stakes. With LDP-fied explanations, we mitigate the privacy breach from explanation (as mentioned in Section 3) and we do not share the prediction vector, but only the label, to protect against MIA attacks from it. However, we acknowledge that our proposed pipeline is inappropriate for systems with a sheer requirement of both prediction vector and useful private explanations. From our findings, it is clear that, at least with DP models, building such systems is not possible. One may argue that if we LDP-fy the prediction vector coming from a non-private model by adding calibrated noise, it could be a better option than discarding the non-private prediction vector. We do not advocate making the prediction vector LDP by injecting noise, as excessive noise could alter the actual prediction and/or mislead decision-makers with false prediction confidence, impairing both model interpretability and utility. We note that sharing the model’s prediction (only the label) is the least possible amount of privacy risk we take to make our pipeline functional. Our findings show that reconciling privacy and explainability in high-stakes situations comes with several other silent factors that need to be exclusively addressed, and is non-trivial. While this study aims to serve as a rigorous starting point, canonically reconciling privacy with explainability, with minimum compromise, is still an open challenge that demands further investigation, probably even beyond the setup and constraints we discussed in this paper.